OpenStudy (anonymous):

how to differentiate y=e^(x/a^2)

OpenStudy (hba):

What is the objective of this question sir?

OpenStudy (anonymous):

how to differentiate

ganeshie8 (ganeshie8):

use chain rule

OpenStudy (anonymous):

just show me it plz

ganeshie8 (ganeshie8):

sure :) corrected typo before : \(\large y=e^{x/a^2}\)

ganeshie8 (ganeshie8):

check ur notes and confirm if it is \(y\) on the lft side....

OpenStudy (anonymous):

no no its x

OpenStudy (anonymous):

ohhh sorry it is y

ganeshie8 (ganeshie8):

cool :)

ganeshie8 (ganeshie8):

\(\large y=e^{x/a^2}\) \(\large \frac{dy}{dx}=\frac{d}{dx}(e^{x/a^2})\)

ganeshie8 (ganeshie8):

wats the derivative of e^x ?

OpenStudy (anonymous):

lnx

ganeshie8 (ganeshie8):

nope. derivative of e^x is just e^x

ganeshie8 (ganeshie8):

ln x is the inverse of e^x, its not the derivative ok ?

ganeshie8 (ganeshie8):

\(\large y=e^{x/a^2}\) \(\large \frac{dy}{dx}=\frac{d}{dx}(e^{x/a^2})\) \(\large \frac{dy}{dx}=e^{x/a^2} * \frac{d}{dx}(x/a^2)\)

OpenStudy (anonymous):

ok

ganeshie8 (ganeshie8):

since we dont have just e^x, we need to take derivative of exponent again. its called chain rule.

ganeshie8 (ganeshie8):

see if above looks okay :)

OpenStudy (anonymous):

yeah but what about d(x/a^2)

ganeshie8 (ganeshie8):

good question :) we still need to work it out

ganeshie8 (ganeshie8):

\(a\) is just a constant. so pull it out

OpenStudy (anonymous):

okay

ganeshie8 (ganeshie8):

\(\large y=e^{x/a^2}\) \(\large \frac{dy}{dx}=\frac{d}{dx}(e^{x/a^2})\) \(\large \frac{dy}{dx}=e^{x/a^2} * \frac{d}{dx}(x/a^2)\) \(\large \frac{dy}{dx}=e^{x/a^2} * \frac{1}{a^2}\frac{d}{dx}(x)\)

ganeshie8 (ganeshie8):

fine ?

ganeshie8 (ganeshie8):

wats the derivative of \(x\) ?

ganeshie8 (ganeshie8):

derivative of x is just 1

ganeshie8 (ganeshie8):

\(\large y=e^{x/a^2}\) \(\large \frac{dy}{dx}=\frac{d}{dx}(e^{x/a^2})\) \(\large \frac{dy}{dx}=e^{x/a^2} * \frac{d}{dx}(x/a^2)\) \(\large \frac{dy}{dx}=e^{x/a^2} * \frac{1}{a^2}\frac{d}{dx}(x)\) \(\large \frac{dy}{dx}=e^{x/a^2} * \frac{1}{a^2} * 1\) \(\large \frac{dy}{dx}=\frac{e^{x/a^2}}{a^2}\)

OpenStudy (anonymous):

okay thank u so much if u have a facebook or eamil to keep in touch

ganeshie8 (ganeshie8):

np... u wlc :) pm me ur email... il add u :)

OpenStudy (anonymous):

brsa999@hotmail.com

ganeshie8 (ganeshie8):

ive copied... delete it here it may be misused by others... lol